If a bacterial messenger RNA that has a coding region of 42 bases long that starts with an AUG and ends with the stop codon UAG. How many different polypeptides are possible from an mRNA?

39

36

12

2012

2014

436
Answer #1

20^12

There are 42 bases.

1 codon = 1 amino acid = 3 bases

# codons = 42/3 = 14

Out of these 14 codons, 1st will be AUG and last will be UAG. We are sure for these two positions so, we will not consider these two. So, now, 12 codons will be left which will be occupied by any one of the 20 amino acids.

# total amino acid = 20

At each of the 12 positions, one out of 20 amino acids can come.

# possible combinations of peptide with 12 codons = 20 × 20 × 20 × 20 × 20 × 20 × 20 × 20 × 20 × 20 × 20 × 20 = 20^12

Please rate high.